OpenStudy (anonymous):

solve for x. log(x)-log(x+9)=1

OpenStudy (anonymous):

Use the fact that log(a) - log(b) = log(a/b) for your first step

OpenStudy (anonymous):

Yeah i got \[\log(x/x+9)=1\] what do i do from here

OpenStudy (anonymous):

Do you know the rule for adding and subtracting logs? Two hundred years ago engineers and scientists made tables of logs of every number because large number multiplications are turned into addition with logs. Say you want to know what 137890*98900 is equal to: If you have a table of logs, you will know that log(137890)+log(98900) = log(137890*98900) = 10.133 and 10^10.133 = 1.358*10^10

OpenStudy (anonymous):

So, from where you are, exponentiate both sides by 10.

OpenStudy (anonymous):

what does that mean ?

OpenStudy (anonymous):

if you know that log(pi) = x, then use each side of the equality as exponents to make a new equation. On one side you'll have \[10^{\log(\pi)}\] and on the other side \[10^{x}\] the \[10^{\log(\pi)}\] simplifies to pi

OpenStudy (jdoe0001):

\(\bf log(x)-log(x+9)=1\implies log\left(\cfrac{x}{x+9}\right)=1\implies log_{10}\left(\cfrac{x}{x+9}\right)=1\\ \quad \\ \textit{log cancellation rule of }\quad a^{log_ax}=x\qquad thus\\ \quad \\ \Large log_{10}\left(\cfrac{x}{x+9}\right)=1\implies 10^{log_{10}\left(\frac{x}{x+9}\right)}=10^1\\ \quad \\\implies \cfrac{x}{x+9}=10\)

OpenStudy (jdoe0001):

hmmm shoot got a bit truncated

OpenStudy (jdoe0001):

\(\bf log(x)-log(x+9)=1\implies log\left(\cfrac{x}{x+9}\right)=1\\ \quad \\\implies log_{10}\left(\cfrac{x}{x+9}\right)=1\\ \quad \\ \textit{log cancellation rule of }\quad a^{log_ax}=x\qquad thus\\ \quad \\ \large log_{10}\left(\cfrac{x}{x+9}\right)=1\implies 10^{log_{10}\left(\frac{x}{x+9}\right)}=10^1\\ \quad \\ \implies \cfrac{x}{x+9}=10\)

OpenStudy (anonymous):

Thank both of you for your help. :)

OpenStudy (anonymous):

Hey @jdoe0001, how do you get your equation fonts to be bold and use the arrow?

OpenStudy (jdoe0001):

\bf <--- turns on BoldFace \implies -> \(\implies\)

OpenStudy (anonymous):

@brendafn happy to help.

OpenStudy (anonymous):

Thanks

OpenStudy (jdoe0001):

@Matt.Mawson http://www.onemathematicalcat.org/MathJaxDocumentation/TeXSyntax.htm

OpenStudy (jdoe0001):

:)

OpenStudy (anonymous):

@jdoe0001 do you know how i would check, because when i plug in 10 to the orginal equation i get -0.27875